OpenStudy (anonymous):

Express A and A^-1 as products of elementary matrices: A=[2 1 1; 1 2 1; 1 1 2].

OpenStudy (cruffo):

we want to find a series of matrices E_1, E_2, ... E_n so that \[E_nE_{n-1} \cdots_2E_1A = I\] So \[A^{-1} = E_nE_{n-1} \cdots_2E_1\] and \[A = E_n^{-1}E_{n-1}^{-1} \cdots E_2^{-1}E_1^{-1}\]

OpenStudy (cruffo):

To find the elementary matrices, row reduce A to echelon form, but encode those row operations as a left matrix multiplication. For instance, the first row operation would be to multiply row 2 by -2 and add to row 1. This would be multipling A on the left by \[\left(\begin{array}{ccc} 1&0&0\\ 1&-2&0\\ 0&0&1\\ \end{array}\right)\]

OpenStudy (cruffo):

which would be E_1

OpenStudy (cruffo):

but, I'll be honest. Seems like we could take advantage of symmetry in this problem.... just don't remember how.

OpenStudy (anonymous):

I don't quite get what you're saying, so have that matrix and column of zeros on the right?

OpenStudy (anonymous):

And try to row reduce that matrix?

OpenStudy (cruffo):

Yes, you want to row reduce A to echelon form, but you want to "capture" each single row operation as a matrix.

OpenStudy (anonymous):

so it'd be like [2 11; 1 2 1; 1 1 2] but everytime I do an operation I'd also have to do the same operation to [1 0 0; 0 1 0; 0 0 1]?

OpenStudy (cruffo):

Yes, by record each change in the identity matrix as it's own matrix, E_1, E_2, ... and make sure you only do one row operation at a time.

OpenStudy (anonymous):

Oh okay thank you!

OpenStudy (cruffo):

\[E_1 = \left(\begin{array}{ccc} 1&0&0\\ 1&-2&0\\ 0&0&1\\ \end{array}\right)\] \[E_2 = \left(\begin{array}{ccc} 1&0&0\\ 0&1&0\\ 1&0&-2\\ \end{array}\right)\]

OpenStudy (cruffo):

\[E_3 = \left(\begin{array}{ccc} 1&0&0\\ 0&1&0\\ 0&1&-3\\ \end{array}\right)\] \[E_4 = \left(\begin{array}{ccc} 1&0&0\\ 0&1&0\\ 0&0&\frac{1}{8}\\ \end{array}\right)\]

OpenStudy (cruffo):

\[E_5 = \left(\begin{array}{ccc} 1&0&0\\ 0&1&1\\ 0&0&1\\ \end{array}\right)\] \[E_6 = \left(\begin{array}{ccc} 1&0&0\\ 0&-\frac{1}{3}&0\\ 0&0&1\\ \end{array}\right)\]

OpenStudy (cruffo):

\[E_7 = \left(\begin{array}{ccc} 1&0&-1\\ 0&1&0\\ 0&0&1\\ \end{array}\right)\] \[E_8 = \left(\begin{array}{ccc} 1&-1&0\\ 0&1&0\\ 0&0&1\\ \end{array}\right)\] \[E_9 = \left(\begin{array}{ccc} \frac{1}{2}&0&0\\ 0&1&0\\ 0&0&1\\ \end{array}\right)\]

OpenStudy (cruffo):

so \[A^{-1} = E_9E_8E_7E_6E_5E_4E_3E_2E_1\]

OpenStudy (anonymous):

nice job :)